the 6th term of an AP is 2 and its common difference is greater than 1. The value of the common difference of the progression so that the product of the 1st,4th and 5th terms is greatest is

Grade:11

1 Answers

Arun
5 years ago
a + 5d = 2
d >1
Now
a1* a4* a5 = (a)(a+ 3d)(a+ 4d) = (2 -5d)(2 -2d)(2 -d)
= 2(2- 5d)(1- d)(2 - d) = f(d)
We have to maximize this
Hence on solving
f'(d) = - (15 d² - 34d +16)
 
Similarly
f''(d) = -(30 d -34)
f(d) is maximum when f'(d) = 0 & f''(d)
On solving
d = 2/3,  8/5
For d = 8/5,  f"(d) = -14
Hence d = 8/5 is the answer